Consider a multinomial experiment involving

n = 150 trials and k = 4 cells. The observed frequencies

resulting from the experiment are shown

in the accompanying table, and the null hypothesis

to be tested is as follows:

H0: p1 = .3, p2 = .3, p3 = .2, p4 = .2

Cell 1 2 3 4

Frequency 38 50 38 24

Test the hypotheses, using  = .05.


1
Expert's answer
2021-01-11T18:55:41-0500

Given a multinomial experiment involving n=150 trials and k=4 cells. For this the data is as follows

Cell 1 2 3 4

Frequency 38 50 38 24

The given null hypothesis is

H0: p1 = 0.3, p2 = 0.3, p3 = 0.2, p4 = 0.2

Alternative hypothesis is

H1: At least one p1 is not equil to its specific value

The test statistic for the Chi-Square goodness-of-fit (χ2) is given by,

"χ^2 = \sum^{k}_{i=1}\frac{(f_i-e_i)^2}{e_i}"

ei = expected frequencies

fi = observed frequencies

Given level of significance α=0.05

Using MINITAB, the following are the steps to get the χ2-statistic for the Account receivable data.

1) Click Stat, Table, and Chi-square Goodness-of-fit Test (One Variable)….

2) Type the observed values into the Observe Counts and specify the variable name.

3) Click Proportions specified by historical counts and Input constants type the values of the proportions under the null hypothesis (0.3, 0.3, 0.2, 0.2)

4) Click OK.

We get the following output



From the above output,

The test statistic χ2= 4.97778

The P-value is 0.173

Here we observe that, the P-value is greater than the level of the significance 0.05, so we fail to reject the null hypothesis. Therefore, there is enough evidence to support the claim that the given proportions are equal to the specified values.